OpenStudy (anonymous):

3 1/8x - 22/5 + 22/5= 7 1/2 3 1/8x= 15/2 + 22/5 25/8x - 25/8= 15/2 + 22/5 25/8x= 3/1 + 11/1 25/8x - 25/8= 14/1- 25/8 x=7/1-25/4 x=18/4 x= 4-1/2 Is this right?

OpenStudy (amistre64):

is x in the denom?

OpenStudy (anonymous):

Hmm good question @amistre. I took it to be as follows: \[\Large \begin{array}{l} {\rm{3 1}}/{\rm{8x }} - {\rm{ 22}}/{\rm{5 }} + {\rm{ 22}}/{\rm{5}} = {\rm{ 7 1}}/{\rm{2}}\\ 3\frac{1}{8}x - \frac{{22}}{5} + \frac{{22}}{5} = 7\frac{1}{2}\\ \frac{{25}}{8}x = \frac{{15}}{2}\\ x = \frac{{\frac{{15}}{2}}}{{\frac{{25}}{8}}}\\ x = \frac{{8 \cdot 15}}{{25 \cdot 2}} = \frac{{120}}{{50}} = \frac{{12}}{5} \end{array}\]

OpenStudy (amistre64):

the wolf pops out x=1/36 if under; and x=36 if not

OpenStudy (anonymous):

I think the Wolf and I are misunderstanding each other. If (3+(1/8))x then: http://www.wolframalpha.com/input/?i=%283%2B%281%29%2F%288%29%29x-%2822%29%2F%285%29%2B%2822%29%2F%285%29%3D7%2B%281%29%2F%282%29